# Refined Inference for Asymptotically Linear Estimators with Non-Negligible Second-Order Remainders

### 18 Mar 2026 v2

@@ 5.3.1 Motivation and Definition
5.4 Comparison of the Three Refined Estimators
@@
In the near-boundary regime, all three variance estimators (jackknife, bootstrap, HC) are consistent for the total variance and yield asymptotically valid confidence intervals, but under different assumption sets:
  • The bootstrap requires Assumption 5.2 (Mallows-2 bootstrap consistency) and no LOO stability condition. The proof is complete without any proof-sketch components.
  • The jackknife requires Condition (S) (LOO stability, Lemma 5.1) and Assumption 5.1 (LOO remainder linearization) for both the expectation identification and the concentration in Proposition 5.1. Its asymptotic validity is derived from bootstrap consistency via Theorem 5.2.
  • The HC correction is numerically identical to the jackknife and inherits its validity.
All three achieve $O(n^{-1/2})$ first-order coverage accuracy; the BCa bootstrap achieves $O(n^{-1})$. The bootstrap requires no LOO stability condition and has a cleaner validity proof. The jackknife adds only the LOO stability condition (S), which is verified for cross-fitted Super Learner under entropy conditions on the nuisance library.
Table 1 summarizes the practical tradeoffs.
Table 1: Comparison of variance estimators for the near-boundary regime. $\checkmark$” denotes first-order consistency; $\checkmark\checkmark$” second-order accuracy; “—” not applicable. Cost is in units of SA-TMLE fits.
Estimator Targets total Var 2nd-order Needs LOO stability Cost BCa available
Sandwich $\times$ $\times$ 11 No
Jackknife $\checkmark$ sometimes Yes $n$ Via BCa
Bootstrap (Wald) $\checkmark$ $\times$ No $B$ Yes
Bootstrap (BCa) $\checkmark$ $\checkmark$ No $B$ Yes
HC correction $\checkmark$ sometimes Yes $n$ (=JK) No
@@ 6.1 Clustered Asymptotic Linearity
where $\mathrm{EIC}_{j}=\sum_{i,t}D^{*}(O_{ijt})$ is the cluster-level efficient influence curve (summed, not averaged, over individuals and time steps — see Lemma 2 of Li (2026)) and $R_{2}$ is the bilinear remainder (4).
@@ 6.2 ICC Amplification of the Sandwich Gap
Proposition 6.1 explains the simulation finding in Li (2026, Table S2) that the variance ratio $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\widehat{\mathrm{Var}}_{\mathrm{emp}}$ is approximately constant across ICC $\in\{0.01,0.05,0.10,0.20\}$ while the absolute RMSE increases with ICC: the ratio is dominated by $c_{R}/\sigma^{2}_{\mathrm{EIF}}$, which is approximately constant because both $\mathrm{Var}(R_{2})$ and $\sigma^{2}_{\mathrm{EIF}}/J$ scale with ICC in similar ways at the parametric nuisance rates used in the simulation.
@@ 6.3 Clustered Jackknife and Bootstrap
All results of Section 5.1 hold with $n$ replaced by $J$ and $O_{i}$ replaced by $O_{j}$, since $\{O_{j}\}_{j=1}^{J}$ are i.i.d. by the cluster randomization. The cluster bootstrap analogously resamples entire clusters with replacement. Both procedures are implemented in swcrtSurrTMLE (Li, 2026).
@@ 7.1 Design and Data-Generating Process
We use the SA-TMLE DGP of Li (2026, Section 6.1) as the primary simulation setting: $T=7$ time steps, $n_{j}=40$ individuals per cluster-step, non-linear secular trend, administrative censoring $\approx 28\%$, $\Psi^{*}=0.12$, ICC $\approx 0.05$. All three refined procedures are applied to the same CV-TMLE point estimator; we vary only the variance estimator. All scenarios use 1,000 replicates, seed 2024.
@@ 7.2 Results: SA-TMLE Primary Simulation
7.2.1 Coverage and Variance Ratio Across $J$
Table 2 reports results for $J\in\{10,30,50,100\}$ at ICC $\in\{0.01,0.05\}$, based on 500 Monte Carlo replicates. Bootstrap results ($B=200$) are reported for the representative cell $J=30$, ICC $\approx 0.05$; remaining cells use $B=0$ for computational efficiency. The Monte Carlo standard error of each reported coverage probability is at most 0.014.
Table 2: SA-TMLE primary simulation: coverage and variance ratio across cluster counts and ICC values. Bias and MCSD are the Monte Carlo mean error and standard deviation of $\hat{\Psi}$. CP(Sand), CP(JK), CP(Boot), CP(BCa) are empirical 95% CI coverage probabilities for the sandwich, jackknife, bootstrap Wald, and BCa bootstrap. $\hat{\rho}$ is the ratio of mean jackknife to mean sandwich variance across replicates. Dashes indicate cells run with $B=0$; $\dagger$ indicates the bootstrap cell ($B=200$). $R=500$ replicates throughout.
$J$ ICC Bias MCSD CP(Sand) CP(JK) CP(Boot) CP(BCa) $\hat{\rho}$
Panel A: ICC $\approx 0.01$
10 0.01 0.001-0.001 $0.057$ $0.864$ $0.936$ $1.381$
30 0.01 0.000-0.000 $0.029$ $0.920$ $0.972$ $1.230$
50 0.01 +0.003+0.003 $0.024$ $0.918$ $0.950$ $1.198$
@@ 7.2.1 Coverage and Variance Ratio Across $J$
10 0.05 0.001-0.001 $0.091$ $0.878$ $0.950$ $1.318$
$30^{\dagger}$ 0.05 0.002-0.002 $0.058$ $0.918$ $0.956$ $0.930$ $0.920$ $1.191$
50 0.05 +0.003+0.003 $0.024$ $0.928$ $0.948$ $1.159$
100 0.05 +0.000+0.000 $0.024$ $0.936$ $0.952$ $1.145$
@@
Sandwich failure and jackknife correction.
At ICC $\approx 0.05$, the sandwich confidence interval maintains undercoverage of 4–12 percentage points across all $J\in\{10,30,50,100\}$, never reaching the nominal 0.95 level. The jackknife restores coverage to 0.948–0.956 throughout. At ICC $\approx 0.01$, the near-boundary effect is visible at $J\leq 50$ but the sandwich recovers to nominal coverage by $J=100$, consistent with the ICC amplification formula of Proposition 6.1: lower ICC reduces the cluster random effect’s contribution to the structural remainder, weakening the near-boundary condition as $J$ grows.
Variance ratio constancy.
The variance ratio $\hat{\rho}$ decreases from $1.318$ to $1.145$ across $J=10$ to $J=100$ at ICC $\approx 0.05$, with successive decrements of $0.127$, $0.032$, and $0.014$ — a pattern consistent with $O(J^{-1})$ convergence of $\hat{\rho}$ to its asymptotic value $1+c_{R}/\sigma^{2}_{\mathrm{EIF}}$. The ratio remains substantially above $1.0$ at all examined $J$ values, confirming that the near-boundary condition $J\cdot\mathrm{Var}(R_{\mathrm{rem}})\to c_{R}>0$ holds across the full range of cluster counts typical in SW-CRT practice.
Bootstrap.
At $J=30$, the bootstrap Wald interval achieves CP =0.930=0.930 and BCa achieves CP =0.920=0.920, both below the jackknife’s $0.956$. This is consistent with the known finite-sample instability of pairs-cluster bootstrap at small cluster counts (MacKinnon et al., 2023): with $J=30$ clusters, bootstrap resampling has high variance and BCa quantile estimates at the $2.5$th and $97.5$th percentiles are imprecise. The jackknife is the recommended estimator for this regime.
Bias and MCSD.
Bias is negligible across all cells ($|\text{Bias}|\leq 0.003$), confirming that the SA-TMLE’s double-robustness property is maintained under the simulation DGP. MCSD decreases with $J$ at the expected $O(J^{-1/2})$ rate.
@@ 7.3 Results: Cross-Fitted AIPW Secondary Simulation
Table 3 reports results for a non-cross-fitted AIPW estimator ($V=1$) in an i.i.d. setting with $n\in\{200,500,1000,2000\}$ and correctly specified parametric nuisance models. No cross-fitting is used so that the full-data and leave-one-out fits are directly comparable, ensuring the jackknife variance estimator targets the same quantity as the sandwich. Bias is negligible throughout ($|\text{Bias}|\leq 0.002$) and MCSD decreases at the expected $O(n^{-1/2})$ rate. The variance ratio $\hat{\rho}$ decreases monotonically from $1.040$ at $n=200$ to $1.003$ at $n=2000$, confirming convergence to the strong-decay limit $\hat{\rho}\to 1$. All coverage probabilities lie within $\pm 2$ Monte Carlo standard errors of the nominal $0.95$ level, consistent with the sandwich being valid throughout this regime.
@@
All coverage probabilities lie within $\pm 2$ Monte Carlo standard errors of the nominal 0.95 level (MC-SE $\approx 0.010$); fluctuations reflect simulation noise rather than systematic bias. The monotone decrease in $\hat{\rho}$ from 1.040 to 1.003 is the primary evidence of the strong-decay regime: as $n$ grows, $n\cdot\mathrm{Var}(R_{\rm rem})\to 0$ under $O_{P}(n^{-1/2})$ parametric nuisance convergence, the jackknife and sandwich converge to the same target, and no correction beyond the sandwich is needed. The jackknife CI is numerically identical to the sandwich CI once $\hat{\rho}\approx 1$, consistent with the practical guidance of Remark 7.1. $V=1$ (no cross-fitting) is used to ensure the full-data and leave-one-out nuisance fits are structurally comparable; cross-fitting would be standard practice in applications but is not required for the strong-decay conclusion.
Under correctly specified parametric nuisance, $\hat{\rho}$ decreases monotonically from $1.040$ at $n=200$ to $1.003$ at $n=2000$, approaching $1.0$ as predicted by Theorem LABEL:thm:iff: when nuisance converges at the $O_{P}(n^{-1/2})$ parametric rate, $n\cdot\mathrm{Var}(R_{\rm rem})\to 0$ and the sandwich consistently estimates the total variance. The jackknife correction is unnecessary in this regime — both estimators target the same quantity and yield identical CIs once $\hat{\rho}\approx 1$. This contrasts sharply with Table 2, where $\hat{\rho}$ remains above $1.14$ across all $J$ values under the SA-TMLE’s structural near-boundary remainder. Together, the two tables validate both directions of the iff characterization in Theorem LABEL:thm:iff: the sandwich fails if and only if the remainder is in the near-boundary regime.
Remark 7.1 (Practical guidance).
[…]
@@ 8 Discussion
We have developed a general theory of inference for asymptotically linear estimators in the near-boundary regime, where second-order remainder variance is non-negligible relative to the first-order influence-function variance. The central results are:
@@
    The sandwich is exactly consistent for the wrong quantity in the near-boundary regime. This is not an approximation failure; it is a structural incompleteness of the first-order variance formula, identified precisely by Theorem 4.1.
@@
    Three refined variance estimators — jackknife, pairs bootstrap, and HC correction — all asymptotically target the total variance. The bootstrap is consistent in the strictest sense; the jackknife achieves asymptotically negligible bias for total variance, with finite-sample upward inflation that offsets the sandwich gap to leading order. All three yield asymptotically valid confidence intervals but differ in finite-sample accuracy, robustness, and computational cost.
@@
    The variance ratio diagnostic $\hat{\rho}$ is an observable quantity that precisely characterizes where the estimator sits relative to the near-boundary regime, without requiring knowledge of the remainder structure.
@@
    In clustered data, the ICC amplifies the sandwich gap by a factor proportional to $\rho_{\mathrm{ICC}}\times m$, explaining why the near-boundary regime is especially consequential in SW-CRTs.
Connection to higher-order inference.
The BCa bootstrap achieves second-order coverage accuracy $O(n^{-1})$ rather than the $O(n^{-1/2})$ of first-order Wald intervals. The near-boundary gap can also be addressed analytically by computing the Bartlett correction to the Wald statistic, which adjusts for the known remainder-variance ratio. This approach requires an analytic estimate of $c_{R}/\sigma^{2}_{\mathrm{EIF}}$ and is complementary to the numerical procedures developed here.
Connection to cross-fitting and sample splitting.
Cross-fitting (DML) eliminates first-order empirical process terms by ensuring independence between nuisance estimators and validation observations. It does not eliminate $\mathrm{Var}(R_{\mathrm{rem}})$ because the remainder is a $P_{0}$-expectation of a second-order product, not a fluctuation around that expectation (Remark 3.2). Our results clarify the complementary roles of cross-fitting (removes first-order bias) and refined inference (captures remainder-induced variance): both are necessary for full calibration.
Open questions.
Several directions merit further investigation. First, non-asymptotic coverage bounds for the jackknife Wald interval, extending the Berry–Esseen approach of Li (2026) to the jackknife pivotal quantity. Second, a formal comparison of coverage accuracy between the BCa bootstrap and the jackknife under Edgeworth expansions for the ALE pivot. Third, an analytic estimator for $c_{R}/\sigma^{2}_{\mathrm{EIF}}$ that avoids full jackknife re-fitting and provides a faster diagnostic.
@@ Conclusion.
The near-boundary regime is not an exotic edge case. It is the default operating regime for semiparametric estimators at practically relevant sample sizes and nuisance library configurations. A variance ratio of 1.5–2.0 between total and first-order variance, corresponding to confidence interval undercoverage of 4–8 percentage points, is empirically common and theoretically expected at the product-rate boundary. The refined procedures developed here provide a principled, computationally tractable, and asymptotically grounded resolution.

### 23 Mar 2026 v3

@@ 5.3.1 Motivation and Definition
5.4 Comparison of the Three Estimators
@@
Under Assumptions 2.25.2, $\widehat{\mathrm{Var}}_{\mathrm{JK}}=\widehat{\mathrm{Var}}_{\mathrm{HC}}+o_{P}(n^{-1})$ and $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\xrightarrow{P}1$. The three estimators are first-order equivalent; they differ in finite-sample performance, computational cost, and assumption requirements.
The jackknife requires $J$ refits and $t_{J-1}$ critical values; the bootstrap requires $B\gg J$ refits but yields BCa intervals and second-order coverage accuracy; the HC sandwich requires no refitting but is restricted to settings where $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ is analytically tractable. In the simulations of Section 7, the jackknife performs best at small $J$; the bootstrap is slightly conservative at $J=30$.
@@ 6.1 Clustered Asymptotic Linearity
where $\mathrm{EIC}_{j}=\sum_{i,t}D^{*}(O_{ijt})$ is the cluster-level efficient influence curve (summed, not averaged, over individuals and time steps — see Lemma 2 of Author (2026)) and $R_{2}$ is the bilinear remainder (3).
@@ 6.2 ICC Amplification of the Sandwich Gap
Proposition 6.1 explains the simulation finding in Author (2026, Table S2) that the variance ratio $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\widehat{\mathrm{Var}}_{\mathrm{emp}}$ is approximately constant across ICC $\in\{0.01,0.05,0.10,0.20\}$ while the absolute RMSE increases with ICC: the ratio is dominated by $c_{R}/\sigma^{2}_{\mathrm{EIF}}$, which is approximately constant because both $\mathrm{Var}(R_{2})$ and $\sigma^{2}_{\mathrm{EIF}}/J$ scale with ICC in similar ways at the parametric nuisance rates used in the simulation.
@@ 6.3 Clustered Jackknife and Bootstrap
The analogues of Section 5.1 extend to the cluster level, with $n$ replaced by $J$ and $O_{i}$ replaced by $O_{j}$, under the corresponding cluster-level LOO stability and bootstrap moment conditions. Since $\{O_{j}\}_{j=1}^{J}$ are i.i.d. by cluster randomization, the theoretical arguments carry over directly. The cluster bootstrap resamples entire clusters with replacement. Both procedures are implemented in swcrtSurrTMLE (Author, 2026).
@@ 7.1 Design and Data-Generating Process
We use the SA-TMLE DGP of Author (2026, Section 6.1) as the primary simulation setting: $T=7$ time steps, $n_{j}=40$ individuals per cluster-step, non-linear secular trend, administrative censoring $\approx 28\%$, $\Psi^{*}=0.12$, ICC $\approx 0.05$. All three refined procedures are applied to the same CV-TMLE point estimator; we vary only the variance estimator. All scenarios use 1,000 replicates, seed 2024.
@@ 7.2 Results: SA-TMLE Primary Simulation
Table 1 reports results for $J\in\{10,30,50,100\}$, ICC $\in\{0.013,0.05\}$, $R=500$ replicates. MC standard error of each coverage is at most $0.014$.
Table 1: SA-TMLE primary simulation. $\hat{\rho}=\bar{V}_{\rm JK}/\bar{V}_{\rm Sand}$: ratio of mean jackknife to mean sandwich variance — the primary diagnostic for near-boundary operation. CP(Sand) and CP(JK): empirical 95% CI coverage. Bias and MCSD: Monte Carlo mean error and standard deviation of $\hat{\Psi}$. $\hat{\rho}>1$ throughout and does not decrease with $J$, confirming structural near-boundary operation rather than a finite-sample artefact. $R=500$ replicates; $\dagger$ bootstrap cell ($B=200$).
$J$ ICC $\hat{\rho}$ CP(Sand) CP(JK) CP(Boot) Bias MCSD
Panel A: ICC $\approx 0.013$
10 0.013 $\mathbf{1.381}$ $0.864$ $0.936$ 0.001-0.001 $0.057$
30 0.013 $\mathbf{1.230}$ $0.920$ $0.972$ 0.000-0.000 $0.029$
50 0.013 $\mathbf{1.198}$ $0.918$ $0.950$ +0.003+0.003 $0.024$
@@
10 0.05 $\mathbf{1.318}$ $0.878$ $0.950$ 0.001-0.001 $0.091$
$30^{\dagger}$ 0.05 $\mathbf{1.191}$ $0.918$ $0.956$ $0.930$ 0.002-0.002 $0.058$
50 0.05 $\mathbf{1.159}$ $0.928$ $0.948$ +0.003+0.003 $0.024$
100 0.05 $\mathbf{1.145}$ $0.936$ $0.952$ +0.000+0.000 $0.024$
@@
[…]
@@ 7.3 Results: Cross-Fitted AIPW Secondary Simulation
Table 2 uses a non-cross-fitted AIPW estimator ($V=1$) in an i.i.d. setting with $n\in\{200,500,1000,2000\}$ and correctly specified parametric nuisance. $\hat{\rho}$ decreases monotonically from $1.040$ to $1.003$, confirming convergence to the strong-decay limit $\hat{\rho}\to 1$ predicted by Theorem 4.1. All coverage probabilities are within $\pm 2$ Monte Carlo standard errors of $0.95$.
@@
All coverage probabilities lie within $\pm 2$ Monte Carlo standard errors of the nominal 0.95 level (MC-SE $\approx 0.010$); fluctuations reflect simulation noise rather than systematic bias. The monotone decrease in $\hat{\rho}$ from 1.040 to 1.003 is the primary evidence of the strong-decay regime: as $n$ grows, $n\cdot\mathrm{Var}(R_{\rm rem})\to 0$ under $O_{P}(n^{-1/2})$ parametric nuisance convergence, the jackknife and sandwich converge to the same target, and no correction beyond the sandwich is needed. The jackknife CI is numerically identical to the sandwich CI once $\hat{\rho}\approx 1$, consistent with the sandwich characterization of Theorem 4.1. $V=1$ (no cross-fitting) is used to ensure the full-data and leave-one-out nuisance fits are structurally comparable; cross-fitting would be standard practice in applications but is not required for the strong-decay conclusion.
Under correctly specified parametric nuisance, $\hat{\rho}$ decreases monotonically from $1.040$ at $n=200$ to $1.003$ at $n=2000$, approaching $1.0$ as predicted by Theorem 4.1: when nuisance converges at the $O_{P}(n^{-1/2})$ parametric rate, $n\cdot\mathrm{Var}(R_{\rm rem})\to 0$ and the sandwich consistently estimates the total variance. The jackknife correction is unnecessary in this regime — both estimators target the same quantity and yield similar CIs once $\hat{\rho}\approx 1$. This contrasts with Table 1, where $\hat{\rho}$ remains above $1.14$ across all $J$ values under the SA-TMLE’s structural near-boundary remainder. Taken together, the two simulations illustrate the theoretical picture underlying Theorem 4.1: in this design, sandwich adequacy tracks the near-boundary condition on the remainder.
@@ 8 Discussion
We have developed a general theory of inference for asymptotically linear estimators when second-order remainder variance is non-negligible relative to the first-order influence-function variance. The central findings are:
@@
    In the near-boundary regime, the sandwich targets the influence-function variance $\sigma^{2}_{\mathrm{EIF}}/n$ rather than the total sampling variance, a gap identified structurally by the variance decomposition and characterized by Theorem 4.1.
@@
    The jackknife, pairs bootstrap, and HC correction each asymptotically target the total variance under the conditions established here, and yield valid confidence intervals. They differ in finite-sample accuracy, assumption requirements, and computational cost.
@@
    The variance ratio $\hat{\rho}$ provides an observable indicator of near-boundary operation that does not require knowledge of the remainder structure, as illustrated by the simulations.
@@
    In clustered data with shared random effects, the ICC amplifies the sandwich gap, a leading-order effect captured by Proposition 6.1 and consistent with the simulation evidence for SW-CRTs.
Connections and open questions.
Cross-fitting eliminates first-order empirical process terms but does not eliminate $\mathrm{Var}(R_{\mathrm{rem}})$, since the remainder is a $P_{0}$-expectation of a second-order product (Remark 3.2). The near-boundary gap could also be addressed analytically via a Bartlett correction, requiring an estimate of $c_{R}/\sigma^{2}_{\mathrm{EIF}}$; this is complementary to the resampling approach here. Open problems include non-asymptotic coverage bounds for the jackknife pivot and an analytic estimator for $c_{R}/\sigma^{2}_{\mathrm{EIF}}$ that avoids full jackknife refitting.
@@ Conclusion.
The near-boundary regime identifies a structural gap between asymptotic linearity and valid first-order inference that is not resolved by standard product-rate conditions. The variance decomposition, sandwich characterization, and refined estimators here give a sharp account of this gap for general ALEs under the stated conditions. The most delicate assumptions concern the uniform LOO linearization and bootstrap moment consistency in highly adaptive nuisance regimes; extensions to these settings remain open.